# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 1.5 Acknowledgement
Most of the results presented in this work were obtained as part of my MSc studies at Ben Gurion University of the Negev, between the years 2019-2021. I would like to thank my MSc advisor (currently my Phd advisor), Eitan Sayag, for teaching me a lot of mathematics and also for many helpful discussions about my research and his many helpful comments during the writing of this text. I would also like to thank Avraham Aizenbud, Joseph Bernstein, Daniel Disegni and Nadya Gurevich. Special thanks to Paul Broussous for pointing out a mistake in a previous version of the proof of Proposition 3.1.
@@ 2 Notations and Preliminaries
In this section we present several well known results that we use in the next sections of this work. These results regard the affine Weyl group, $W_{aff}$ of $G$, the affine building $\mathcal{B}_{G}$, parahoric group schemes, the Iwahori Hecke algebra $\mathcal{H}(G,I)$ and $p$-adic symmetric spaces.
2.1 The affine Weyl group of $G$
@@
Let $T$ be a maximal split torus of $G$, we denote by $T^{0}$ the unique maximal compact subgroup of $T$, if $T\cong(F^{\times})^{n}$ then $T^{0}\cong(\mathcal{O}^{\times})^{n}$.
We denote by $W(T)=N_{G}(T)/T$ the Weyl group of $G$ and $T$, or just by $W$ if it is clear from the context with respect to which maximal torus the Weyl group is taken. Similarly we denote by $W_{aff}(T)=N_{G}(T)/T^{0}$ or $W_{aff}$ the affine Weyl group of $G$ and $T$.
Let $(X^{*}(T),\Phi,X_{*}(T),\Phi^{\vee})$ be the root datum (see [Spr98] for the definition) that correspond to $T$. We will use the notation $X_{*}=X_{*}(T)$ and $X^{*}=X^{*}(T)$.
@@
The group $W$ is isomorphic to the Coxeter group generated by reflections over $\Delta$ (see [Bor91]).
For any root $\alpha\in\Phi$ denote by $\alpha^{\vee}\in\Phi^{\vee}$ the coroot of $\alpha$ (see [Spr98] for a definition).
@@
There is a perfect pairing between $X_{*}\times X^{*}\rightarrow Aut(\mathbb{G}_{m})=\mathbb{Z}$ given by composition.
Let $V=X_{*}\otimes\mathbb{R}$, for any $\alpha\in\Phi$ let $s_{\alpha}\in N_{G}(T)$ be an element that induces the reflection over $\alpha^{\vee}$ in $V$ (see [Lan95] lemma 0.19). We consider $\alpha\in\Phi$ as an element of $V^{*}$. For any integer $k\in\mathbb{Z}$ we consider the affine root $\alpha+k$ and the following affine reflection on V $x\rightarrow s_{\alpha}(x)-k\alpha^{\vee}$.
By proposition 2.3 each affine root $\alpha+k$ induces a function on V.
@@
Our definition of chambers is slightly different then the standard one. In the standard definition, the chambers are open subsets of $V$.
Let $\mathcal{C}$ be the unique chamber on which all the simple roots are positive and which contains 00. $\mathcal{C}$ is called the fundamental alcove.
We define a length function $l:W_{aff}(T)\rightarrow\mathbb{N}$, for $w\in W_{aff}(T)$ define $l(w)$ to be the number of walls between $\mathcal{C}$ and $w\mathcal{C}$.
The group $W$ is generated by $\{s_{\alpha}|\alpha\in\Delta\}$ (see [D.B] theorem 6) and it acts on $V$ normalizing the lattice $\mathbb{Z}\Phi^{\vee}$. Let $W_{a}=\mathbb{Z}\Phi^{\vee}\rtimes W\subset Aff(V)$.
@@
The group $W_{a}$ is a Coxeter group generated by reflections over the faces of $\mathcal{C}$. $l$ restricted to $W_{a}$ is the equal to the Coxeter length function on $W_{a}$. Additionally $W_{aff}(T)=W_{a}\rtimes\Omega$, where $\Omega=l^{-1}(0)\cong X^{\vee}/\mathbb{Z}\Phi^{\vee}$.
Denote by $\tilde{\Delta}$ a set of elements of $N_{G}(T)$ which induce on $V$ the reflections over the faces of the fundamental alcove, $\mathcal{C}$.
@@ 2.2 The affine building of $G$
There are several ways to construct the affine building of a split p-adic group $G$, we follow the construction presented in [DS11] (for more details see [Lan95]).
@@
Let $\mathbf{T}\subset\mathbf{G}$ be a maximal torus, let $\Phi$ be its root system. For a root $\alpha\in\Phi$ there is a unipotent group $\mathbf{U}_{\alpha}\subset G$ and an isomorphism $u_{\alpha}:\mathbb{G}_{a}\rightarrow\mathbf{U}_{\alpha}$ such that for any $t\in T$, $t^{-1}u_{\alpha}(a)t=u_{\alpha}(\alpha(t)a)$ (see [Bor91] proposition 21.9).
@@
There is a filtration on $\mathbb{G}_{a}(F)$ given by powers of the uniformizer times $\mathcal{O}$, this induces via $u_{\alpha}$ a filtration on $U_{\alpha}$. Denote $U_{\alpha,k}=u_{\alpha}(\pi^{k}\mathcal{O})$ for $k\in\mathbb{Z}$. This yields a filtration on $U_{\alpha}$.
@@
We define $N_{L}$ to be the group of elements in $N_{G}(T)$ that fixes $L$ pointwise.
@@
We denote by $P_{L}$ the group generated by $U_{L}$ and $N_{L}$.
For a single point $x$ we denote $U_{x}=U_{\{x\}}$ and $P_{x}=P_{\{x\}}$.
Now we can give a description of the enlarged affine building $\mathcal{B}_{G}$ of the $p$-adic group $G$.
@@
All maximal split tori of $G$ are conjugate (see Theorem 18.14 in [Mil17]).
@@
$g^{-1}T_{1}g=T_{2}$. Conjugation by $g$ induces an isomorphism of $G$ topological spaces between $\mathcal{B}^{T_{1}}_{G}$ and $\mathcal{B}^{T_{2}}_{G}$. (See [Lan95] proposition 9.20).
@@
All apartments have a structure of an affine space which they carry from $V$.
@@
$\mathcal{B}_{G}$ is an abstract poly-simplicial complex that satisfies the following properties:
@@
    Every $n-1$ poly-simplex in an apartment $\mathcal{A}$ lies in exactly two adjacent $n$ poly-simplices of $\mathcal{A}$ and the graph of adjacent $n$ poly-simplices is connected.
@@
    If two poly-simplices both lie in apartments $\mathcal{A}$ and $\mathcal{A}^{\prime}$, then there is a simplicial isomorphism of $\mathcal{A}$ onto $\mathcal{A}^{\prime}$ fixing the vertices of the two simplices. This isomorphism can be taken to be action by an element of $G$.
(For more details see section 9 of [Lan95]).
The poly-simplices of dimension $n$ in $\mathcal{B}_{G}$ are called chambers, they are all of the from $g\mathcal{C}$. Chambers of $\mathcal{B}^{nen}_{G}$ are quotients of chambers of $\mathcal{B}_{G}$ with respect to $\mathcal{Z}$.
The group of element in $G$ that fix a specific chamber in $\mathcal{B}_{G}$ pointwise is called an Iwahori subgroup of $G$.
The next theorem summarizes some important properties of the building mentioned in subsections 1.11 and 1.13 in [DS11].
@@
    For any $L\subset\mathcal{A}$, the group $P_{L}$ is exactly the group of elements that fix $L$ pointwise.
@@
Theorem 2.14 (The Affine Bruhat Decomposition, [IM65]).
@@
Let $\mathcal{C}$ be a chamber of $\mathcal{B}_{G}$, let $I$ be the Iwahori subgroup which stabilizes $\mathcal{C}$ pointwise and let $J$ be the group that stabilizes $\mathcal{C}$ (not necessarily pointwise).
Using theorems 2.14 and 2.5 we can construct a group homomorhpism $\omega:G\rightarrow\Omega$ by $\omega(i_{1}xwi_{2})=w$ for $i_{1},i_{2}\in I$, $x\in W_{a}$ and $w\in\Omega$.
@@
    $I$ is normal in $J$ and the homomorphism $\omega$ induces an isomorphism $J/I\cong\Omega$.
Let $I$ be an Iwahori subgroup of $G$, theorem 2.15 allows us to think about $I\backslash G$ as the set of $\Omega$ colored chambers in the building of $G$. Denote by $C_{G}$ the set of chambers in $\mathcal{B}_{G}$, we denote the set of $\Omega$ colored chambers by $C_{\Omega}(G)=C_{G}\times\Omega$. $G$ acts on $C_{\Omega}(G)$ diagonally, when the action on $\Omega$ is given by the left action via $\omega:G\rightarrow\Omega$, meaning $g(\mathcal{C},o)=(g\mathcal{C},\omega(g)o)$.
Two more important results which we will use in this work are the convexity of apartments and the Iwahori decomposition.
@@
The convexity of apartments has two forms, the first is given by the following theorem (See [Lan95] Proposition 9.6).
@@
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be two chambers, a path of adjacent chambers which starts at $\mathcal{C}_{1}$ and ends in $\mathcal{C}_{2}$ is called a gallery. It will be referred to as a minimal path or a minimal gallery if the number of chambers it contains is equal to the minimal number needed to pass from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ via adjacent chambers.
@@
If $\mathcal{C}_{1},\mathcal{C}_{2}$ are chambers in an apartment $\mathcal{A}$ then any minimal path between them is contained in $\mathcal{A}$. ([Bro89] chapter 4 section 4).
@@
This is a generalization of the classical Iwahori decomposition. In the case of $G=GL_{n}(F)$ the statement of the classical Iwahori decomposition is the following:
Let $K=GL_{n}(\mathcal{O})$, $I\subset K$ be the subgroup of matrices reduced to upper-triangular when taken mod $p$. Let $B\subset GL_{n}(F)$ be the Borel subgroup of upper-triangular matrices and $B^{t}$ the Borel subgroup of lower triangular matrices. Let $T$ be the maximal split torus of diagonal matrices. Let $U$ be the unipotent of $B$ and $U^{t}$ the unipotent of $B^{t}$. Denote $I^{+}=I\cap U$ and $I^{-}=I\cap U^{t}$. The classical Iwahori decompisiton states that
@@
This can be recovered from Theorem 2.19 by taking $L=\mathcal{C}$ the fundamental alcove. Then we will have $P_{L}=I$, $U^{+}_{L}=I^{+}$ and $U^{-}_{L}=I^{-}$.
@@
Theorem 2.21 (The building construction is a map of groupoids, [BT84] §4.2.12).
@@ 2.3 Parahoric group schemes
In this subsection we describe certain group schemes defined over $\mathcal{O}$, related to the Burhat Tits building, we follow [Tit79], [McN10], [Lan95] and [KP23]. In this subsection we will refer to subsets of $\mathcal{B}_{G}$ which are bounded mod center, meaning their image in $\mathcal{B}_{G}^{nen}$ is bounded, as bounded subsets.
Let $E$ be an unramified extension of $F$ and let $\mathcal{B}^{E}_{G}$ be the enlarged affine building of $\mathbf{G}(E)$. Here we denote $\mathcal{B}_{G}$ by $\mathcal{B}^{F}_{G}$. There is a canonical embedding $\mathcal{B}^{F}_{G}\subset\mathcal{B}^{E}_{G}$, let $L\subset\mathcal{B}^{F}_{G}$ be a bounded set. We denote by $P^{E}_{L}$ the point-wise stabilizer of $L$ under the action of $\mathbf{G}(E)$ on $\mathcal{B}_{G}^{E}$.
@@
Let $L$ be a bounded set which is contained in some apartment of $\mathcal{B}^{F}_{G}$. There exists a smooth affine group scheme ${\cal G}_{L}$, defined over $\mathcal{O}$, unique up to a unique isomorphism such that:
@@
    For every unramified Galois extension $E/F$ with ring of integers $\mathcal{O}_{E}$ we have ${\cal G}_{L}(\mathcal{O}_{E})\cong P^{E}_{L}$.
(See 3.4.1 in [Tit79]).
In our case, where $G$ is split over $F$, the group ${\cal G}_{L}$ can be constructed explicitly, as done in subsection 6.2 of [McN10].
@@
Let $\alpha+k$ be an affine root of $T$ where $\alpha$ is a root of $T$ and $k\in\mathbb{Z}$. Let $u_{\alpha}:\mathbb{G}_{a}\rightarrow\mathbf{U}_{\alpha}$ be the isomorphism from proposition 2.5. Let $\pi$ be a uniformeizer of $\mathcal{O}$ and let $\mathbb{G}^{k}_{a}$ be the smooth $\mathcal{O}$ scheme whose $\mathcal{O}$ points are $\pi^{k}\mathcal{O}$. Let ${\cal U}_{\alpha+k}$ be the smooth group scheme obtained from $\mathbb{G}^{k}_{a}$ and $u_{\alpha}$ such that its $\mathcal{O}$ points are $u_{\alpha}(\pi^{k}\mathcal{O})$.
@@
Let $L$ be a bounded subset of $\mathcal{A}$. The inclusion of $T^{0}$ and $U_{\alpha_{L}}$ into $P_{L}$ can be extended to isomorphisms from $\mathbf{T}$ and ${\cal U}_{\alpha_{L}}$ onto closed subschmes of ${\cal G}_{L}$.
Like in theorem 2.19 we can define ${\cal U}^{+}_{L}=\Pi_{\alpha\in\Phi^{+}}{\cal U}_{\alpha_{L}}$ and similarly ${\cal U}^{-}_{L}$. The product of the maps from ${\cal U}_{\alpha_{L}}$ to ${\cal G}_{L}$ yields an isomorphism from ${\cal U}^{-}_{L}$ and ${\cal U}^{+}_{L}$ onto closed subschemes of ${\cal G}_{L}$.
Lastly, The product ${\cal U}^{+}_{L}\times\mathbf{T}\times{\cal U}^{-}_{L}$ maps isomorphiclly onto an open subscheme of ${\cal G}_{L}$.
For the proof see 6.1 in [Lan95].
Denote by $\overline{{\cal G}}_{L}$ the special fiber of ${\cal G}_{L}$. $\overline{{\cal G}}_{L}$ is an algebraic group over the residue field $k$.
@@
Let $\Phi_{C}$ be the set of roots constant on $\mathcal{C}$. The root system of $\overline{{\cal G}}_{L}/R_{u}(\overline{{\cal G}}_{L})$ is $\Phi_{C}$. (See 8.4.10 in [KP23])
@@
$\overline{{\cal G}}_{L}$ has a unique Levi factor which contains $\mathbf{T}(k)$. Here a Levi factor means a direct complement to the unipotent radical.
Proof.
This is stated without a proof in [Tit79] and it is theorem A of [McN10]. ∎
@@
Let $L_{1}\subset L_{2}$ two bounded subsets of some apartment, the identity map $G\rightarrow G$ extends in a unique way to a $\mathcal{O}$ group homomorphism $res^{L_{2}}_{L_{1}}:{\cal G}_{L_{1}}\rightarrow{\cal G}_{L_{2}}$. (See 6.2 in [Lan95]).
@@
Let $L$ be any bounded subset of an apartment and let $f$ be a face of some chamber such that $f\cap L\neq\varnothing$ and $f$ is maximal with this property. Then $res^{L_{2}}_{L_{1}}$ is an isomorphism on the reductive quotients of the special fibers. (See 3.5 in [Tit79]).
Let $L_{1}\subset L_{2}$ be two faces of some chamber in the building, the map $res^{L_{2}}_{L_{1}}$ induces a map $\overline{res}^{L_{2}}_{L_{1}}$ from $\overline{{\cal G}}_{L_{2}}$ to $\overline{{\cal G}}_{L_{1}}/R_{u}(\overline{{\cal G}}_{L_{1}})$.
@@
Let $L$ be a face of some chamber in the building, denote by $X_{L}$ the poset of all faces of any chamber which contain $L$, the ordering is with respect to inclusion. The map
@@
is an order preserving bijection between $X_{L}$ and the spherical building of $\overline{{\cal G}}_{L}/R_{u}(\overline{{\cal G}}_{L})$ (See 9.22 in [Lan95]).

### v3

@@ 1.5 Acknowledgement
Most of the results presented in this work were obtained as part of my MSc studies at Ben-Gurion University of the Negev. I would like to thank my MSc advisor (currently my PhD advisor), Eitan Sayag, for teaching me a lot of mathematics and also for many helpful discussions about my research and his many helpful comments during the writing of this text. I would also like to thank Avraham Aizenbud, Joseph Bernstein, Daniel Disegni and Nadya Gurevich. Special thanks to Paul Broussous for pointing out a mistake in a previous version of the proof of Proposition 3.1. I would like to thank the referees for their careful reading and their helpful comments and suggestions. I was partially supported by ISF grant No. 1781/23 during the work on this paper.
@@ 2 Notations and Preliminaries
In this section we present several well known results that we use in this work. These results regard the extended affine Weyl group, $W_{aff}$ of $G$, the affine building $\mathcal{B}_{G}$, parahoric group schemes, the Iwahori Hecke algebra $\mathcal{H}(G,I)$ and $p$-adic symmetric spaces.
2.1 The extended affine Weyl group of $G$
@@
Let $T$ be a maximal split torus of $G$, we denote by $T^{0}$ the unique maximal compact subgroup of $T$, if $T\cong(F^{\times})^{n}$ then $T^{0}\cong(\mathcal{O}^{\times})^{n}$, here $\mathcal{O}$ is the integer ring of $F$.
We denote by $W(T)=N_{G}(T)/T$ the Weyl group of $G$ and $T$, or just by $W$ if it is clear from the context with respect to which maximal torus the Weyl group is taken. Similarly, we denote by $W_{aff}(T)=N_{G}(T)/T^{0}$ or just by $W_{aff}$ the extended affine Weyl group of $G$ and $T$.
Let $(X^{*}(T),\Phi,X_{*}(T),\Phi^{\vee})$ be the root datum (see [32] for the definition) that corresponds to $T$.
We will use the notation $X_{*}=X_{*}(T)$ and $X^{*}=X^{*}(T)$.
@@
The group $W$ is isomorphic to the Coxeter group generated by reflections over $\Delta$ (see [5]).
For any root $\alpha\in\Phi$ denote by $\alpha^{\vee}\in\Phi^{\vee}$ the coroot of $\alpha$ (see [32] for a definition).
@@
There is a perfect pairing $X_{*}\times X^{*}\rightarrow Aut(\mathbb{G}_{m})=\mathbb{Z}$ given by composition.
Let $V=X_{*}\otimes\mathbb{R}$, for any $\alpha\in\Phi$ let $s_{\alpha}\in N_{G}(T)$ be an element that induces the reflection over $\alpha^{\vee}$ in $V$ (see [21] Lemma 0.19). We consider $\alpha\in\Phi$ as an element of $V^{*}$. For any integer $k\in\mathbb{Z}$ we consider the affine root $\alpha+k$ and the following affine reflection on $V$ defined by $x\mapsto s_{\alpha}(x)-k\alpha^{\vee}$.
By Proposition 2.3 each root $\alpha$ induces a function on $V$. Each affine root $\alpha+k$ also induces a function on $V$.
@@
Our definition of chambers is slightly different than the standard one. In the standard definition, the chambers are open subsets of $V$.
Let $\mathcal{C}$ be the unique chamber on which all the simple roots are positive and which contains 0. $\mathcal{C}$ is called the fundamental chamber.
We define a length function $l:W_{aff}(T)\rightarrow\mathbb{N}$. For $w\in W_{aff}(T)$ define $l(w)$ to be the number of walls between $\mathcal{C}$ and $w\mathcal{C}$.
The group $W$ is generated by $\{s_{\alpha}|\alpha\in\Delta\}$ (see [32] Theorem 7.1.9) and it acts on $V$ preserving the lattice $\mathbb{Z}\Phi^{\vee}$. Let $W_{a}=\mathbb{Z}\Phi^{\vee}\rtimes W$, It is a subgroup of affine transformations of $V$.
@@
The group $W_{a}$ is a Coxeter group generated by reflections over the facets of $\mathcal{C}$. The function $l$ restricted to $W_{a}$ is the equal to the Coxeter length function on $W_{a}$. Additionally $W_{aff}(T)=W_{a}\rtimes\Omega$, where $\Omega=l^{-1}(0)\cong X^{\vee}/\mathbb{Z}\Phi^{\vee}$.
Let $\tilde{\Delta}$ be a set of elements in $N_{G}(T)$ which induce on $V$ the reflections over the facets of the fundamental chamber $\mathcal{C}$.
@@ 2.2 The affine building of $G$
There are several ways to construct the affine building of a split $p$-adic group $G$, we follow the construction presented in [12] (for more details see [21]).
@@
Let $\mathbf{T}\subset\mathbf{G}$ be a maximal torus. Let $\Phi$ be its root system. For a root $\alpha\in\Phi$ there is a unipotent group $\mathbf{U}_{\alpha}\subset G$ and an isomorphism $u_{\alpha}:\mathbb{G}_{a}\rightarrow\mathbf{U}_{\alpha}$ such that for any $t\in T$, $t^{-1}u_{\alpha}(a)t=u_{\alpha}(\alpha(t)a)$ (see [5] Proposition 21.9).
@@
There is a filtration on $\mathbb{G}_{a}(F)$ given by powers of the uniformizer times $\mathcal{O}$, this induces via $u_{\alpha}$ a filtration on $U_{\alpha}$. Let $\pi$ be a uniformizer and denote $U_{\alpha,k}=u_{\alpha}(\pi^{k}\mathcal{O})$ for $k\in\mathbb{Z}$. This yields a filtration on $U_{\alpha}$.
@@
We define $N^{1}_{L}$ to be the group of elements in $N_{G}(T)$ that fixes $L$ pointwise.
@@
We denote by $P^{1}_{L}$ the group generated by $U_{L}$ and $N^{1}_{L}$.
For a single point $v\in V$ we denote $U_{v}=U_{\{v\}}$ and $P_{v}=P_{\{v\}}$.
Now we can give a description of the enlarged (also called extended) affine building $\mathcal{B}_{G}$ of the $p$-adic group $G$.
@@
All maximal split tori of $G$ are conjugate (see Theorem 18.14 in [25]).
@@
$g^{-1}T_{1}g=T_{2}$. Conjugation by $g$ induces an isomorphism of $G$ topological spaces between $\mathcal{B}^{T_{1}}_{G}$ and $\mathcal{B}^{T_{2}}_{G}$. (See [21] Proposition 9.20).
@@
All apartments have a structure of an affine space.
The chambers of the enlarged building are defined to be the sets of the form $g\mathcal{C}=\{(g,v)|v\in\mathcal{C}\}/\sim$. Chambers of $\mathcal{B}^{nen}_{G}$ are quotients of chambers of $\mathcal{B}_{G}$ with respect to $\mathcal{Z}$.
@@
$\mathcal{B}_{G}$ is a poly-simplicial complex that satisfies the following properties:
@@
    Every n1n-1 poly-simplex in an apartment $\mathcal{A}$ lies in exactly two adjacent $n$ poly-simplices of $\mathcal{A}$ and the graph of adjacent $n$ poly-simplices is connected.
@@
    If two poly-simplices both lie in apartments $\mathcal{A}$ and $\mathcal{A}^{\prime}$, then there is a simplicial isomorphism of $\mathcal{A}$ onto $\mathcal{A}^{\prime}$ fixing the poly-simplices. This isomorphism can be taken to be an action by an element of $G$.
(For more details see Section 9 of [21]).
The poly-simplices of dimension $n$ in $\mathcal{B}_{G}$ are the chambers of $\mathcal{B}_{G}$.
The group of elements in $G$ that fix a specific chamber in $\mathcal{B}_{G}$ pointwise is called an Iwahori subgroup of $G$ (see Definition 4.1.3 in [19]).
The next theorem summarizes some important properties of the building mentioned in subsections 1.11 and 1.13 in [12].
@@
    For any $L\subset\mathcal{A}$, the group $P^{1}_{L}$ is exactly the group of elements that fix $L$ pointwise.
@@
Theorem 2.14 (The Affine Bruhat Decomposition, [18]).
@@
Let $\mathcal{C}$ be a chamber of $\mathcal{B}_{G}$, let $I$ be the Iwahori subgroup which fixes $\mathcal{C}$ pointwise and let $J$ be the group that fixes $\mathcal{C}$ as a set (not necessarily pointwise).
Using Theorems 2.14 and 2.5 we can construct a group homomorphism $\omega:G\rightarrow\Omega$ by $\omega(i_{1}xoi_{2})=o$ for $i_{1},i_{2}\in I$, $x\in W_{a}$ and $o\in\Omega$.
@@
    II is normal in $J$ and the homomorphism $\omega$ induces an isomorphism $J/I\cong\Omega$.
Definition 2.16.
For any subset $L$ inside an apartment $\mathcal{A}$, we define $P_{L}=P^{1}_{L}\cap Ker(\omega)$. This is the group of elements in $Ker(\omega)$ that fix $L$ pointwise.
Let $I$ be an Iwahori subgroup of $G$. Proposition 2.15 allows us to think about $G/I$ as the set of $\Omega$ colored chambers in the building of $G$. Denote by $C_{G}$ the set of chambers in $\mathcal{B}_{G}$, we denote the set of $\Omega$ colored chambers by $C_{\Omega}(G)=C_{G}\times\Omega$. The group $G$ acts on $C_{\Omega}(G)$ diagonally, when the action on $\Omega$ is given by $\omega:G\rightarrow\Omega$, meaning $g(\mathcal{C},o)=(g\mathcal{C},\omega(g)o)$.
@@
The convexity of apartments has two forms, the first is given by the following theorem (See [21] Proposition 9.6).
@@
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be two chambers, a path of adjacent chambers which starts at $\mathcal{C}_{1}$ and ends at $\mathcal{C}_{2}$ is called a gallery. It will be referred to as a minimal path or a minimal gallery if the number of chambers in it is equal to the minimal number needed to pass from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ via adjacent chambers.
@@
If $\mathcal{C}_{1},\mathcal{C}_{2}$ are chambers in an apartment $\mathcal{A}$ then any minimal gallery between them is contained in $\mathcal{A}$. (See [8] Chapter 4 Section 4).
@@
This is a generalization of the classical Iwahori decomposition. In the case of $G=GL_{n}(F)$ the statement of the classical Iwahori decomposition is as follows:
Let $K=GL_{n}(\mathcal{O})$, $I\subset K$ be the subgroup of matrices reduced to upper-triangular when taken mod $\pi$. Let $B\subset GL_{n}(F)$ be the Borel subgroup of upper-triangular matrices and $B^{t}$ the Borel subgroup of lower triangular matrices. Let $T$ be the maximal split torus of diagonal matrices. Let $U$ be the unipotent of $B$ and $U^{t}$ the unipotent of $B^{t}$. Denote $I^{+}=I\cap U$ and $I^{-}=I\cap U^{t}$. The classical Iwahori decompisiton states that
@@
This can be recovered from Theorem 2.20 by taking $L=\mathcal{C}$ the fundamental chamber. Then we will have $P_{L}=I$, $U^{+}_{L}=I^{+}$ and $U^{-}_{L}=I^{-}$.
@@
Theorem 2.22 (The building construction is a map of groupoids, [9] §4.2.12).
@@ 2.3 Parahoric group schemes
In this subsection we describe certain group schemes defined over $\mathcal{O}$, related to the Burhat Tits building, we follow [33], [24], [21] and [19]. In this subsection we will refer to subsets of $\mathcal{B}_{G}$ which are bounded mod center, meaning their image in $\mathcal{B}_{G}^{nen}$ is bounded, as bounded subsets.
Let $\mathcal{O}$ be the integer ring of $F$.
@@
Let $L$ be a bounded set which is contained in some apartment of $\mathcal{B}^{F}_{G}$. There exists a unique smooth affine group scheme ${\mathcal{G}}_{L}$, defined over $\mathcal{O}$ such that:
@@
    We have ${\mathcal{G}}_{L}(\mathcal{O})\cong P_{L}$.
(See Section 8.3 in [19]. The notations in [19] are $G(k)^{0}_{\Omega}$ for the groups $P_{L}$, and $\mathscr{G}^{0}_{\Omega}$ for ${\mathcal{G}}_{L}$).
In our case, where $G$ is split over $F$, the group ${\mathcal{G}}_{L}$ can be constructed explicitly, as done in Subsection 6.2 of [24].
@@
Let $\alpha+k$ be an affine root of $T$ where $\alpha$ is a root of $T$ and $k\in\mathbb{Z}$. Let $u_{\alpha}:\mathbb{G}_{a}\rightarrow\mathbf{U}_{\alpha}$ be the isomorphism from Proposition 2.6. Let $\pi$ be a uniformizer of $\mathcal{O}$ and let $\mathbb{G}^{k}_{a}$ be the smooth $\mathcal{O}$ scheme whose $\mathcal{O}$ points are $\pi^{k}\mathcal{O}$. Let ${\mathcal{U}}_{\alpha+k}$ be the smooth group scheme obtained from $\mathbb{G}^{k}_{a}$ and $u_{\alpha}$ such that its $\mathcal{O}$ points are $u_{\alpha}(\pi^{k}\mathcal{O})$.
@@
Let $L$ be a bounded subset of $\mathcal{A}$. The inclusion of $T^{0}$ and $U_{\alpha_{L}}$ into $P_{L}$ can be extended to isomorphisms from $\mathbf{T}$ and ${\mathcal{U}}_{\alpha_{L}}$ onto closed sub-schemes of ${\mathcal{G}}_{L}$.
Like in Theorem 2.20 we can define ${\mathcal{U}}^{+}_{L}=\Pi_{\alpha\in\Phi^{+}}{\mathcal{U}}_{\alpha_{L}}$ and similarly ${\mathcal{U}}^{-}_{L}$. The product of the maps from ${\mathcal{U}}_{\alpha_{L}}$ to ${\mathcal{G}}_{L}$ yield an isomorphism from ${\mathcal{U}}^{-}_{L}$ and ${\mathcal{U}}^{+}_{L}$ onto closed sub-schemes of ${\mathcal{G}}_{L}$.
Lastly, The product ${\mathcal{U}}^{+}_{L}\times\mathbf{T}\times{\mathcal{U}}^{-}_{L}$ maps isomorphiclly onto an open sub-scheme of ${\mathcal{G}}_{L}$.
For the proof see either 6.1 in [21] or 8.3.14 in [19].
Denote by $\overline{{\mathcal{G}}}_{L}$ the special fiber of ${\mathcal{G}}_{L}$. $\overline{{\mathcal{G}}}_{L}$ is an algebraic group over the residue field $k$.
@@
Let $\Phi_{L}$ be the set of roots constant on $L$. The root system of $\overline{{\mathcal{G}}}_{L}/R_{u}(\overline{{\mathcal{G}}}_{L})$ is $\Phi_{L}$. (See 8.4.10 in [19])
The following is stated without a proof in [33] and is Theorem A of [24].
@@
$\overline{{\mathcal{G}}}_{L}$ has a unique Levi factor which contains $\mathbf{T}(k)$. Here a Levi factor means a direct complement to the unipotent radical.
@@
Let $L_{1}\subset L_{2}$ be two bounded subsets of some apartment, the identity map $G\rightarrow G$ extends in a unique way to a $\mathcal{O}$ group homomorphism $res^{L_{2}}_{L_{1}}:{\mathcal{G}}_{L_{2}}\rightarrow{\mathcal{G}}_{L_{1}}$. (See either 6.2 in [21] or 8.3.17 in [19]).
@@
Let $L$ be any bounded subset of an apartment and let $f$ be a facet of some chamber such that $f\subset L$ and $f$ is maximal with this property. Assume that $f\subset L$, then $res^{L}_{f}$ is an isomorphism on the reductive quotients of the special fibers. (See 3.5 in [33]).
Let $L_{1}\subset L_{2}$ be two facets of some chamber in the building, the map $res^{L_{2}}_{L_{1}}$ induces a map $\overline{res}^{L_{2}}_{L_{1}}$ from $\overline{{\mathcal{G}}}_{L_{2}}$ to $\overline{{\mathcal{G}}}_{L_{1}}/R_{u}(\overline{{\mathcal{G}}}_{L_{1}})$.
@@
Let $L$ be a facet of some chamber in the building, denote by $X_{L}$ the poset of all facets of any chamber which contain $L$, the ordering is with respect to inclusion. The map
@@
is an order preserving bijection between $X_{L}$ and the spherical building of $\overline{{\mathcal{G}}}_{L}/R_{u}(\overline{{\mathcal{G}}}_{L})$ (See 9.22 in [21]).